OpenStudy (anonymous):

What is the vertex of the graph of y = -1(x - 2)2 + 3? answer choices: (3, -2) (-2, 3) (3, 2) (2, 3)

OpenStudy (anonymous):

....

OpenStudy (anonymous):

y=-1(x-2)^2 +3 y-3=-1(x-2)^2 (x-2)^2=-(y-3) it is a downward parabola with vertex (2,3) ...(by substituting x-2=0 and y-3=0

OpenStudy (anonymous):

so is it 2,3 or -2,-3?

OpenStudy (anonymous):

it is 2,3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw